OpenStudy (anonymous):

when flipping 4 coins, what is the probability of "more heads than tails"?

OpenStudy (anonymous):

satellite73 help me after this one

OpenStudy (anonymous):

that means either 3 heads and one tail or 4 heads the probability of all the possibilities can be seen by looking at the fourth row of pascal's triangle \[1,4,6,4,1\] the denominator of each of the probabilities is \(2^4=16\) so the probabilities are \[\frac{1}{16}, \frac{4}{16},\frac{6}{16},\frac{4}{16}, \frac{1}{16}\] for no heads, one head, two heads, three heads, four heads respectively

OpenStudy (anonymous):

more heads than tails means add up the last two, i.e. \[\frac{4}{16}+\frac{1}{16}\]

OpenStudy (anonymous):

ok i get it now, thx!

OpenStudy (anonymous):

yw